OpenStudy (anonymous):

Solve for x using logarithms. Give an exact answer. Do not give a decimal approximation. 37 = 550e−0.2x

OpenStudy (shane_b):

First, is this the problem?\[\large 37=550e^{-.2x}\]

OpenStudy (anonymous):

yes that is the problem

OpenStudy (shane_b):

\[37=550e^{-.2x}\]First divide both sides by 550 to get:\[\frac{37}{550}=e^{-.2x}\]Then take the natural log of both sides:\[ln(\frac{37}{550})=ln(e^{-.2x})\]\[ln(\frac{37}{550})=(-.2x)lne\] lne=1 so that becomes: \[ln(\frac{37}{550})=-.2x\]Finally, divide both sides by -.2 to isolate x:\[x=\frac{ln(\frac{37}{550})}{-.2}\]

OpenStudy (anonymous):

that's what I put originally and it was marked wrong

OpenStudy (anonymous):

wait, just kidding. I put -0.02 in the denominator the first time..thanks

OpenStudy (shane_b):

np